The pulley is suspended by a cord C. Suppose that m1 = 1.3 kg and m2 = 4.0 kg .
Determine the tension in the cord that supports the pulley C after the masses are released and before one hits the ground. Ignore the mass of the pulley and cords.
1
Expert's answer
2019-02-19T14:29:07-0500

T=m_1 g+m_2 g=(m_1+m_2 )g

T=(1.3+4.0)9.8=52 N.